Three times a number added to 3 is 30. Find the number

Answer by Ms. U(2)

Let x be the number.
3(x)+3= 30 (Transpose 3 to the right side.)
3x =30-3
3x = 27 (Divide both side by 3 to get the value of x.)
3x/3 = 27/3
x = 9